OpenStudy (anonymous):

3!= a. 24 b 9 c 6 d none of the above

OpenStudy (anonymous):

help

OpenStudy (anonymous):

thank you can you help me with more

OpenStudy (yemenrip):

yup

OpenStudy (anonymous):

ill hlp :)

OpenStudy (anonymous):

10!/6!=

OpenStudy (anonymous):

@yemenRIP No direct answers that is against openstudy code of conduct

OpenStudy (anonymous):

10!/6!= 5040

OpenStudy (anonymous):

divide by 2 10 and 6

OpenStudy (anonymous):

5/3=1 2/3

OpenStudy (anonymous):

@iigotek00 No direct answers, what part about that do you guys not understand?

OpenStudy (anonymous):

a 24 b 11,628 c 5,040 d a number to big to compute

OpenStudy (anonymous):

you guy now the answer

OpenStudy (anonymous):

I am here to help, which is something you aren't here to do. So by your logic, you should leave.

OpenStudy (anonymous):

10!/6!= a 24 b 11,628 c 5,040 d a number to big to compute

OpenStudy (anonymous):

Well \[\large \sf \frac{10!}{6!}=\frac{10 \times 9 \times 8 \times 7 \times 6 \times 5 \times 4\times 3 \times 2 \times 1}{6 \times 5 \times 4 \times 3 \times 2 \times 1}\] so just cancel out and then multiply

OpenStudy (anonymous):

WOW good job, yemenRIP gets his gold start today. You sure are smart!!!!!1111!!

OpenStudy (yemenrip):

thanks i guess

OpenStudy (anonymous):

can you help

OpenStudy (yemenrip):

yes

OpenStudy (anonymous):

10!/6!= a 24 b 11,628 c 5,040 d a number to big to compute what is it

OpenStudy (yemenrip):

any more